OpenStudy (anonymous):

(3/x + 2/x+2) / (3/x+2 - 2/x)

OpenStudy (anonymous):

5x + 6 / x-4 thats what i got

OpenStudy (callisto):

\[\frac{\frac{3}{x} + \frac{2}{x+2} }{\frac{3}{x+2} -\frac{2}{x}} = \frac{\frac{3(x+2)+2(x)}{x(x+2)}}{\frac{3(x) - 2(x+2)}{x(x+2)} } = \frac{3x+6+2x}{3x-2x-4} = \frac{5x+6}{x-4}\]

OpenStudy (anonymous):

Yep thanks :)

sam (.sam.):

\[\Huge \frac{\frac{3}{x}+\frac{2}{x+2}}{\frac{3}{x+2}-\frac{2}{x}}\] Multiply each term by a factor of 1 that will equate all the denominators. here, all terms need a denominator of x(x+2). The \(\frac{3}{x}\) expression needs to be multiplied by ((x+2))/((x+2)) to make the denominator x(x+2). \[\Huge \frac{\frac{3\times (x+2)}{x \times (x+2)}+\frac{2}{x+2}\times \frac{x}{x}}{\frac{3}{x+2}-\frac{2}{x}}\] Simplifying top \[\Huge \frac{\frac{5x+6}{x(x+2)}}{\frac{3}{x+2}-\frac{2}{x}}\]